Constants Part A The electric field midway between two equal but opposite point charges is 336 N/C, and the distance between the charges is 11.0 cm What is the magnitude of the charge on each? Express your answer to three significant figures and include the appropriate units. QlValue Units Submit Previous Answers Request Answer
Answer #1

Solution :-

Given that

The electric field mid way between two equal and opposite charge is (E) =336N/C

The distance between the two charge is (r) =11cm =0.11m

We know that equation for the electric field is given by E1 =kq/r2

Then at midway the two electricfields is given by

E =|E1+E2 |=kq/(r/2)2+kq/(r/2)2 =2kq/(r/2)2

Then by simplyfying for magnitude of each charge on each is given by

q =Er2/8k =(336)(0.11)2/8*(9*109) = 36.96×10-9/72=0.5133nC or 5.133*10-10C